Determine whether the given relation is reflexive, symmetric, transitive, or none of these. R is the ”greater than or equal to” relation on the set of real numbers: For all x, y ∈ R , xRy ⇐⇒ x ≥ y 


1
Expert's answer
2021-08-17T08:40:10-0400

Since for any "x \in R\,\,x \ge x" , then "\forall x\,\,xRx" and the relation is reflexive.

Let the conditions "xRy" and "yRx" then "x \ge y" and "y \ge x" . This is only possible if "x=y" . But then the condition "\forall x,y\,\,\,xRy \Rightarrow yRx" is not met. So, relation isn't symmetric.

Let "xRy" and "yRz" . Then "x \ge y" and "y \ge z" . But then "x \ge z" . So, "\forall x,y,z\,\,xRy \wedge yRz \Rightarrow xRz" and relation is transitive.

Answer: relation is reflexive, relation isn't symmetric, relation is transitive.